OpenStudy (anonymous):

point-slope the equation y-6=3(x-2)

OpenStudy (anonymous):

y-6=3(x-2) y-6=3x-6 y=3x-6+6 y=3x since y=mx+c where m is the slope therefore m=3

OpenStudy (anonymous):

y+4=-2(x+1)